OpenStudy (anonymous):

What is the standard form the a circle centered at the region that passes through the point (-2,5)?

OpenStudy (btaylor):

what is the center point?

OpenStudy (anonymous):

Origin

OpenStudy (btaylor):

Thanks. The formula is: \[(y-k)^2 + (x-h)^2 = r^2\] Where (h,k) is the center point and r is the radius. From this, we know that h=0 and k=0 because the center is at (0,0)  We know the radius by the Pythagorean theorem as: \[\sqrt{2^2+5^2} = \sqrt{4+25}=\sqrt{29}\] Therefore, the equation is: \[x^2 + y^2 = 29\]

OpenStudy (anonymous):

Thanks a ton!

OpenStudy (btaylor):

you're welcome.